OpenStudy (anonymous):

Find the slope of the line whose equation is 5y = x - 3. -3 -3/5 1/5 5 -3 ??

OpenStudy (anonymous):

y=ax+b a=slope y=1/5x-3/5 slope=1/5

OpenStudy (anonymous):

Express your equation in the form of slope-intercept form, y = mx + c. From there, the answer is m as it is the slope